OpenStudy (anonymous):

What are the steps in finding the velocity and acceleration functions when x(t)=6 Then how do I find the velocity and acceleration at 2.5s?

OpenStudy (turingtest):

x(t)=6 is your position function? velocity is the derivative of position and acceleration is the derivative of velocity, so v(t)=x'(t)=0 a(t)=v'(t)=x''(t)=0 so a(2.5)=0... how boring... in other words the object in question is just sitting at point x=6, so it is neither accelerating or moving at all. If your position function depended on time t then it might be more interesting. Like if it was x(t)=6t^3+3t+7 v(t)=18t^2+3 a(t)=36t then a(2.5)=36(2.5)=90m/s^2 does that help?

OpenStudy (anonymous):

yes! so that would mean if x(t)=5t^2+17t then v(t)=10t^4+17 a(t)=40t so then a(2.5)=40(2.5)=100m/s^2

OpenStudy (turingtest):

you got the derivatives a wrong, but you have the right idea x(t)=5t^2+17t v(t)=10t+17 a(t)=10 so in this case a(2.5)=10 (t doen't matter if it's not in the equation, right?) remember differentiation: Der(x^n)=nx^n-1 don't know how you got t^4 for your velocity equation

OpenStudy (anonymous):

5-1

OpenStudy (anonymous):

but clearly that was wrong

OpenStudy (turingtest):

subtract 1 from the EXPONENT when you take the derivative. Leave the coefficient alone. Derivative of\[ax^n\]\[=anx ^{n-1}\] I think you need a bit of practice with derivatives. http://tutorial.math.lamar.edu/

OpenStudy (anonymous):

i really do

OpenStudy (turingtest):

no calculus yet?

OpenStudy (anonymous):

nope in calc 2 but i travled last year and skipped pre-calc. so trying to catch up. also in pysics 2 which is going fast than my calc class

OpenStudy (anonymous):

just calc not calc 2...opps

OpenStudy (turingtest):

you'll have derivatives like those down in no time, but it sucks when the physics teachers aren't in sync with the math you're working on. keep at it. :)

OpenStudy (anonymous):

thanks!